### v1

# On Limiting Probability Distributions of Higher Order Markov Chains

Abstract
The limiting (or stationary) probability distribution is one of the key characteristics of a Markov chain since it shows its long-term behavior. In this paper, for a higher order Markov chain, we establish a sufficient condition for the existence of its limiting probability distribution. This condition is built upon the regularity of its transition tensor. Our results extend the corresponding conclusions for first order chains. Besides, they complement the existing results concerning higher order chains that rely on approximation schemes or two-phase power iterations.
Keywords: higher order Markov chains, limiting probability distributions, stationary probability distributions, transition tensors, regularity
1 Introduction
Let $X=\{X_{t}:t=1,2,\ldots\}$ be a stochastic process, where $X_{t}$ is a random variable representing the state of a system at time $t$. Assume that $X$ takes values in the state space $S=\{1,2,\ldots,n\}$. Then, $X$ is called a first order Markov chain if the following Markovian property holds:
$\Pr(X_{t+1}=i|X_{t}=j,X_{t-1}=k_{1},\ldots,X_{1}=k_{t-1})=\Pr(X_{t+1}=i|X_{t}=j)$
for any $t\geq 1$ and $i,j,k_{1},\ldots,k_{t-1}\in S$. This property can be interpreted as that the future evolution of the process $X_{t+1}$ depends on its current status $X_{t}$ only.
We also assume that the above chain $X$ is homogeneous, meaning that $\Pr(X_{t+1}=i|X_{t}=j)$ is independent of $t$. Under this assumption, we denote
$\Pr(X_{t+1}=i|X_{t}=j)=p_{ij},$
where $i,j\in S$. These are the transition probabilities of the chain. Accordingly, $P=[p_{ij}]\in\mathbb{R}^{n\times n}$ is the transition matrix, which is column stochastic, namely $0\leq p_{ij}\leq 1$ for any $i,j\in S$ and $\sum_{i=1}^{n}p_{ij}=1$ for any $j\in S$.
The preceding Markovian property can be adapted in a way so that the future evolution of the process depends not only on its current status but also on one or more of its immediately past statuses. Formally, $X$ is called an $(m-1)$th order, where $m\geq 2$, Markov chain if the Markovian property takes the following $(m-1)$th order form:
$\begin{array}[]{l}\Pr(X_{t+1}=i_{1}|X_{t}=i_{2},\ldots,X_{t-m+2}=i_{m},\ldots,X_{1}=i_{t+1})\\[5.0pt] =\Pr(X_{t+1}=i_{1}|X_{t}=i_{2},\ldots,X_{t-m+2}=i_{m})\end{array}$ (1.1)
for any $t\geq m-1$ and $i_{1},\ldots,i_{m},\ldots,i_{t+1}\in S$.
Again, under the homogeneity assumption, we denote
$\Pr(X_{t+1}=i_{1}|X_{t}=i_{2},\ldots,X_{t-m+2}=i_{m})=p_{i_{1}i_{2}\ldots i_{m}},$
where $t\geq m-1$ and $i_{1},i_{2},\ldots,i_{m}\in S$, and call it the transition probability from states $(i_{2},\ldots,i_{m})$ to state $i_{1}$. Moreover, the resulting $m$th order, $n$-dimensional tensor ${\cal P}=[p_{i_{1}i_{2}\ldots i_{m}}]$ is the transition tensor, which is “column” stochastic, namely $0\leq p_{i_{1}i_{2}\ldots i_{m}}\leq 1$ for any $i_{1},i_{2},\ldots,i_{m}\in S$ and $\sum_{i_{1}=1}^{n}p_{i_{1}i_{2}\ldots i_{m}}=1$ for any $i_{2},\ldots,i_{m}\in S$. By the term higher order chains, we refer to the case when $m\geq 3$. Clearly, when $m=2$, the chain reduces to a first order one.
For the past decade or so, higher order Markov chains have drawn much attention in the literature, see [1, 2, 3, 4, 7, 8, 9, 10, 11, 12, 13, 20, 21, 26]. A focal point in a majority of these existing works is the limiting probability distribution of a higher order chain. The following definitions, however, includes the first order case as well since it is also relevant here.
DEFINITION
Assume $m\geq 2$. For an $(m-1)$th order (homogeneous) Markov chain $X=\{X_{t}:t=1,2,\ldots\}$ on state space $S=\{1,2,\ldots,n\}$,
$x_{t}=[\Pr(X_{t}=1)~{}\Pr(X_{t}=2)~{}\ldots~{}\Pr(X_{t}=n)]^{T}\in\mathbb{R}^{n}$ (1.2)
is called the probability distribution of $X$ at time $t$.
DEFINITION
Under the assumptions of Definition 1, if $\lim_{t\rightarrow\infty}x_{t}$ exists and, moreover, it does not depend on the initial probability distribution of $X_{1},X_{2},\ldots,X_{m-1}$, then
$\pi=\lim_{t\rightarrow\infty}x_{t}$
is called the limiting probability distribution of the chain $X$.
We observe that $\pi$ itself is a probability distribution, namely $\pi\geq 0$ and $\sum_{i=1}^{n}\pi_{i}=1$. It is an important characteristic of the chain since it provides useful information regarding its long-term behavior. It should be mentioned that $\lim_{t\rightarrow\infty}x_{t}$ may not exist or may depend on the initial probability distribution of $X_{1},X_{2},\ldots,X_{m-1}$. An example about this shall be provided later after necessary background material is introduced.
For a first order chain with transition matrix $P$, a well-known sufficient condition regarding the existence of the limiting probability distribution is $P$ being regular. Denote $P^{k}=[p^{(k)}_{ij}]\in\mathbb{R}^{n\times n}$, where $k=1,2,\ldots$. Then, $p^{(k)}_{ij}$ represents the $k$-step transition probability from states $j$ to $i$, that is,
$p^{(k)}_{ij}=\Pr(X_{t+k}=i|X_{t}=j)$ (1.3)
for any $t\geq 1$. The transition matrix $P$ is called regular if there exists some $k\geq 1$ such that $P^{k}>0$. Under such a condition, $\lim_{t\rightarrow\infty}x_{t}=\pi$ exists and is independent of the initial probability distribution of $X_{1}$, and, in addition, $\pi>0$. Moreover, $\lim_{k\rightarrow\infty}P^{k}=\pi e^{T}$, where $e\in\mathbb{R}^{n}$ is the vector of all ones. For these existing results concerning regular first order chains, we refer the reader to [14, 17].
It cannot be overemphasized that regardless of the order of the chain, its limiting probability distribution, if it exists, does not depend on where the chain starts. In the first order case, this limiting probability distribution is also called the stationary probability distribution. The latter term has been used interchangeably in the literature on higher order chains. We, however, choose the former since for a higher order chain, this distribution lacks the interpretation of stationarity when it comes to the evolution of the chain.
In light of the aforementioned results on first order chains, it is natural to raise the question of whether similar results can be established on higher order chains. As far as we can see, this question has thus far been partially answered by certain approximations of the limiting probability distribution, a two-phase procedure under a stronger regularity requirement, or a matrix framework that only deals with second order chains. In this paper, we shall approach the question here largely from a tensor formulation perspective to provide a more complete answer.
Before proceeding, it should be pointed out that due to consideration of computational tractability, a vast majority of existing works on the limiting probability distribution of a higher order chain have taken the so-called $Z$-eigenvector approach [21] — see a brief survey in Section 3. This approach, however, provides only an approximation of the limiting probability distribution [5, 26]. Moreover, such an approximation, in general, does not reflect the true nature of the limiting probability distribution [26]. In comparison, our results here concern the exact limiting probability distribution.
2 Tensor Formulation and Matricization
The concept of higher order Markov chains dates back to as early as 1930s [16]. Most, if not all, classical references [14, 15], however, only talk briefly about their connection with first order chains but do not provide detailed discussion. In the recent works we have cited in the introductory section, a tensor framework has emerged as the main instrument for studying higher order chains. Nevertheless, results regarding such a framework appear to be a bit scattered, particularly relevant to the limiting probability distribution problem. In this section, therefore, we first set the stage methodically.
Assuming $m\geq 3$, we continue with an $(m-1)$th order (homogeneous) Markov chain $X=\{X_{t}:t=1,2,\ldots\}$ on state space $S=\{1,2,\ldots,n\}$, with an $m$th order, $n$-dimensional transition tensor ${\cal P}=[p_{i_{1}i_{2}\ldots i_{m}}]$.
Similar to the first order case, the $k$-step, with $k\geq 1$, transition probability from states $(i_{2},\ldots,i_{m})$ to state $i_{1}$ is given by
$p^{(k)}_{i_{1}i_{2}\ldots i_{m}}=\Pr(X_{t+k}=i_{1}|X_{t}=i_{2},\ldots,X_{t-m+2}=i_{m})$
for any $t\geq m-1$. In particular, $p^{(1)}_{i_{1}\ldots i_{m}}=p_{i_{1}\ldots i_{m}}$. By convention, we define
$p^{(0)}_{i_{1}i_{2}\ldots i_{m}}=\delta_{i_{1}i_{2}\ldots i_{m}}=\left\{\begin{array}[]{cl}1,&i_{1}=i_{2};\\ 0,&i_{1}\neq i_{2}.\end{array}\right.$
These $k$-step transition probabilities form a stochastic tensor too, which is denoted by ${\cal P}^{(k)}=[p^{(k)}_{i_{1}\ldots i_{m}}]$ and has the same size as $\cal P$. The special case of $k=0$ leads to the so-called identity tensor ${\cal I}=[\delta_{i_{1}i_{2}\ldots i_{m}}]$ — note that this identity tensor is different from those in [18, 23]. In addition, it is known that the $k$-step transition probabilities can be determined by the following recurrence formula [1, 9]:
$p^{(k+1)}_{i_{1}i_{2}\ldots i_{m}}=\sum_{j=1}^{n}p_{ji_{2}\ldots i_{m}}p^{(k)}_{i_{1}ji_{2}\ldots i_{m-1}},~{}k=0,1,2,\ldots.$ (2.1)
In [10], (2.1) is written as a tensor product ${\cal P}^{(k+1)}={\cal P}\boxtimes{\cal P}^{(k)}$. In this sense, ${\cal P}^{(k)}$ can be thought of as the $k$th power of $\cal P$. The identity tensor, on the other hand, satisfies ${\cal P}\boxtimes{\cal I}={\cal P}$. The operation $\boxtimes$ is distributive, yet it is neither commutative nor associative.
Next, we recall the concepts of irreducibility, ergodicity, and regularity when it comes to the higher order chain $X$. These can be extended to nonnegative tensors too, but we shall restrict ourself to the stochastic scenario here. Specifically, the transition tensor $\cal P$ is said to be :
  • irreducible if for each $J$ such that $\emptyset\neq J\subsetneq S$, $p_{i_{1}i_{2}\ldots i_{m}}>0$ for some $i_{1}\in J$ and some $i_{2},\ldots,i_{m}\in J^{c}$.
  • ergodic if given any $i_{1},i_{2},\ldots,i_{m}\in S$, there exists $k\geq 1$, which may depend on $i_{1},i_{2},\ldots,i_{m}$, such that $p^{(k)}_{i_{1}i_{2}\ldots i_{m}}>0$.
  • regular if there exists some $k\geq 1$ such that $p^{(k)}_{i_{1}i_{2}\ldots i_{m}}>0$ for every $i_{1},i_{2},\ldots,i_{m}\in S$.
Obviously, regularity implies ergodicity. In the meantime, it is also known that ergodicity implies irreducibility but the reverse does not hold [9]. The latter is one of the differences between the higher order and the first order cases since irreducibility and ergodicity are the same on a first order chain.
We now review a systematical way the higher order chain $X$ can be reduced to a first order one. Denote $T=\{i_{1}i_{2}\ldots i_{m-1}:i_{1},i_{2},\ldots,i_{m-1}\in S\}$. It is the collection of multi-indices of length $m-1$. These multi-indices are arranged via linear indexing [22]. When $m=4$ and $n=2$, for example,
$T=\{111,211,121,221,112,212,122,222\}.$
Clearly, the cardinality of $T$ is $N=n^{m-1}$. Consider random variables $Y_{t}=(X_{t},X_{t-1},\ldots,X_{t-m+2})$ that take values in $T$. When
$X_{t}=i_{1},X_{t-1}=i_{2},\ldots,X_{t-m+2}=i_{m-1},$
for example, we have $Y_{t}=i_{1}i_{2}\ldots i_{m-1}$. Let $Y=\{Y_{t}:t=m-1,m,\ldots\}$ be the first order chain on state space $T$. It is also referred to as the reduced first order chain obtained from $X$. The transition matrix $Q\in\mathbb{R}^{N\times N}$ of $Y$ is given by
$q_{i_{1}i_{2}\ldots i_{m-1},j_{2}j_{3}\ldots j_{m}}=\Pr(Y_{t+1}=i_{1}i_{2}\ldots i_{m-1}|Y_{t}=j_{2}j_{3}\ldots j_{m}).$
We observe that $q_{i_{1}i_{2}\ldots i_{m-1},j_{2}j_{3}\ldots j_{m}}=p_{i_{1}i_{2}\ldots i_{m-1}j_{m}}$ when $i_{\ell}=j_{\ell}$ for all $\ell=2,3,\ldots,m-1$; otherwise, $q_{i_{1}i_{2}\ldots i_{m-1},j_{2}j_{3}\ldots j_{m}}=0$. Again, considering $m=4$ and $n=2$, we arrive at
$Q=\left[\begin{array}[]{cccccccc}p_{1111}&0&0&0&p_{1112}&0&0&0\\ p_{2111}&0&0&0&p_{2112}&0&0&0\\ 0&p_{1211}&0&0&0&p_{1212}&0&0\\ 0&p_{2211}&0&0&0&p_{2212}&0&0\\ 0&0&p_{1121}&0&0&0&p_{1122}&0\\ 0&0&p_{2121}&0&0&0&p_{2122}&0\\ 0&0&0&p_{1221}&0&0&0&p_{1222}\\ 0&0&0&p_{2221}&0&0&0&p_{2222}\end{array}\right].$
In general, $p_{i_{1}i_{2}\ldots i_{m}}$ is located in $Q$ on row
$i_{1}+n(i_{2}-1)+\ldots+n^{m-2}(i_{m-1}-1)$
and column
$i_{2}+n(i_{3}-1)+\ldots+n^{m-2}(i_{m}-1).$
We shall give an easier way of constructing $Q$ later in this section.
As in (1.3), the $k$-step transition probability of $Y$ from states $j_{2}j_{3}\ldots j_{m}$ to $i_{1}i_{2}\ldots i_{m-1}$ is given by
$\begin{array}[]{rcl}q^{(k)}_{i_{1}i_{2}\ldots i_{m-1},j_{2}j_{3}\ldots j_{m}}&=&\Pr(X_{t+k}=i_{1},X_{t+k-1}=i_{2},\ldots,X_{t+k-m+2}=i_{m-1}\\[8.00003pt] &&\hskip 30.00005pt|X_{t}=j_{2},X_{t-1}=j_{3},\ldots,X_{t-m+2}=j_{m}).\end{array}$ (2.2)
Clearly, for $21\leq k\leq m-$ 2, $q^{(k)}_{i_{1}i_{2}\ldots i_{m-1},j_{2}j_{3}\ldots j_{m}}=0$ unless $i_{k+1}=j_{2},i_{k+2}=j_{3},\ldots,i_{m-1}=j_{m-k}$.
It should be pointed out that problems regarding the higher order chain $X$ cannot always be fully addressed by resorting to the reduced first order chain $Y$. To explain this, we mention several examples here: (1) $Y$ may be non-ergodic even though $X$ is ergodic [10], (2) the mean first passage times of $X$ may not be obtained in a direct fashion from those of $Y$ [10], and (3) a recurrent state in $X$ may no longer be a recurrent state in $Y$ [15]. The reduced first order chain, nevertheless, can still be a useful tool as can be seen shortly.
Moving on, let $P\in\mathbb{R}^{n\times N}$ be the mode-$1$ matricization [19] of the transition tensor $\cal P$. Simply put, $P$ consists of the frontal slices of $\cal P$ which are arranged side by side and ordered according to linear indexing. Taking the case when $m=4$ and $n=2$, $P$ has the form
$P=\left[\begin{array}[]{cc|cc|cc|cc}p_{1111}&p_{1211}&p_{1121}&p_{1221}&p_{1112}&p_{1212}&p_{1122}&p_{1222}\\ p_{2111}&p_{2211}&p_{2121}&p_{2221}&p_{2112}&p_{2212}&p_{2122}&p_{2222}\end{array}\right].$
We can refer to the entries of $P$ using multi-indices in the form of $p_{i_{1},i_{2}\ldots i_{m}}$, which is the same as $p_{i_{1}i_{2}\ldots i_{m}}$. In general, $p_{i_{1}i_{2}\ldots i_{m}}$ is located in $P$ on row $i_{1}$ and column
$i_{2}+n(i_{3}-1)+\ldots+n^{m-2}(i_{m}-1).$
Similarly, $P^{(k)}$ stands for the mode-$1$ matricization of ${\cal P}^{(k)}$. By convention, $P^{(0)}$ is the mode-$1$ matricization of the identity tensor $\cal I$, that is,
$P^{(0)}=[\underbrace{I_{n}~{}I_{n}~{}\ldots~{}I_{n}}_{n^{m-2}}],$ (2.3)
where $I_{n}\in\mathbb{R}^{n\times n}$ is the identity matrix.
Incidentally, the transition matrix $Q$ of the first order chain $Y$ is constructed in [26] in a quite complicated way by using the product of a block diagonal matrix and a permutation matrix. By the mode-$1$ matricization of ${\cal P}^{(k)}$, the construction of $Q$ can be simplified as
$Q=G\ast P,$ (2.4)
where
$G=[\underbrace{I_{n^{m-2}}~{}I_{n^{m-2}}~{}\ldots~{}I_{n^{m-2}}}_{n}]$
and $\ast$ stands for the columnwise Khatri-Rao product [24].
The above mode-$1$ matricization of $\cal P$ can bridge the probability distribution of the higher order chain $X$ to that of the reduced first order chain $Y$. Specifically, let $y_{t}=x_{t,t-1,\ldots,t-m+2}\in\mathbb{R}^{N}$ be the probability distribution of $Y$ at time $t$, whose $i_{1}i_{2}\ldots i_{m-1}$th entry is given by
$\Pr(Y_{t}=i_{1}i_{2}\ldots i_{m-1})=\Pr(X_{t}=i_{1},X_{t-1}=i_{2},\ldots,X_{t-m+2}=i_{m-1})$
for any $i_{1}i_{2}\ldots i_{m-1}\in T$. As usual, these entries are ordered again by linear indexing. As an example, with $m=4$ and $n=2$, we obtain
$y_{t}=\left[\begin{array}[]{c}\Pr(X_{t}=1,X_{t-1}=1,X_{t-2}=1)\\ \Pr(X_{t}=2,X_{t-1}=1,X_{t-2}=1)\\ \Pr(X_{t}=1,X_{t-1}=2,X_{t-2}=1)\\ \Pr(X_{t}=2,X_{t-1}=2,X_{t-2}=1)\\ \Pr(X_{t}=1,X_{t-1}=1,X_{t-2}=2)\\ \Pr(X_{t}=2,X_{t-1}=1,X_{t-2}=2)\\ \Pr(X_{t}=1,X_{t-1}=2,X_{t-2}=2)\\ \Pr(X_{t}=2,X_{t-1}=2,X_{t-2}=2)\end{array}\right].$
By $x_{t}$ in (1.2) and $y_{t}$ as above, it can be shown [21] that
$x_{t+1}=Py_{t}.$ (2.5)
Formula (2.5) is exactly what the two-phase procedure in [26] is based on. Besides, both $x_{t}$ and $y_{t}$ are further formulated as tensors in [26].
With the background material we have presented thus far, it is time to give an example to illustrate that, in general, $\lim_{t\rightarrow\infty}x_{t}$ may depend on the initial probability distribution of $X_{1},X_{2},\ldots,X_{m-1}$.
EXAMPLE 2.1
Let us take a second order, i.e. $m=3$, chain $X$ on state space $\{1,2\}$ whose transition tensor $\cal P$ is as follows:
${\cal P}(:,:,1)=\left[\begin{array}[]{cc}1&1/2\\ 0&1/2\end{array}\right],~{}{\cal P}(:,:,2)=\left[\begin{array}[]{cc}1/2&0\\ 1/2&1\end{array}\right].$
If we start with $X_{1}=1$ and $X_{2}=1$, then $y_{t}=[1~{}0~{}0~{}0]^{T}$ for any $t\geq 1$ and so $\lim_{t\rightarrow\infty}x_{t}=[1~{}0]^{T}$ by (2.5). On the other hand, if we start with $X_{1}=2$ and $X_{2}=2$, then $y_{t}=[0~{}0~{}0~{}1]^{T}$ for any $t\geq 1$ and so $\lim_{t\rightarrow\infty}x_{t}=[0~{}1]^{T}$.
For a second order chain, the following technical lemma can be seen in [25]. We now extend it to the general higher order chain $X$.
LEMMA 2.1
Let $X$ be an $(m-1)$th order Markov chain on state space $S$ whose transition tensor is $\cal P$. Assume $u,v=0,1,2,\ldots$. Let $P^{(u)}$ and $Q$ be the mode-$1$ matricization of ${\cal P}^{(u)}$ and the transition matrix of the reduced first order chain $Y$, respectively. Then,
$P^{(u+v)}=P^{(u)}Q^{v}.$ (2.6)
In particular, we have
$P^{(v)}=P^{(0)}Q^{v}$ (2.7)
for any $v=0,1,2,\ldots$.
Proof: Clearly, (2.6) holds when $u\geq 0$ and $v=0$. Assume now $u=0$ but $v\geq 1$. Note that on row $i_{1}$, $P^{(0)}$ has ones at, in multiindex form, columns $i_{1}j_{2}\ldots j_{m-1}$ for any $j_{2},\ldots,j_{m-1}\in S$ and zeros elsewhere. Hence,
$\begin{array}[]{rcl}(P^{(0)}Q^{v})_{i_{1},i_{2}\ldots i_{m}}&=&\displaystyle\sum_{j_{1}\ldots j_{m-1}}p^{(0)}_{i_{1},j_{1}\ldots j_{m-1}}q^{(v)}_{j_{1}\ldots j_{m-1},i_{2}\ldots i_{m}}\\[16.00008pt] &=&\displaystyle\sum_{j_{2}\ldots j_{m-1}}q^{(v)}_{i_{1}j_{2}\ldots j_{m-1},i_{2}\ldots i_{m}}\\[16.00008pt] &=&\displaystyle\sum_{j_{2}\ldots j_{m-1}}\Pr(X_{t+v}=i_{1},X_{t+v-1}=j_{2},\ldots,X_{t+v-m+2}=j_{m-1}\\ &&\hskip 60.00009pt|X_{t}=i_{2},X_{t-1}=i_{3},\ldots,X_{t-m+2}=i_{m})\\[8.00003pt] &=&p^{(v)}_{i_{1},i_{2}\ldots i_{m}},\end{array}$
where, according to the remark below (2.2), the third summation above is effectively over $j_{2}\ldots j_{v}$ only when $22\leq v\leq m-$ 2 and is equal to $p_{i_{1},i_{2}\ldots i_{m}}$ directly when $v=1$.
Similarly, for $u,v\geq 1$, we have
$\begin{array}[]{l}(P^{(u)}Q^{v})_{i_{1},i_{2}\ldots i_{m}}\\[8.00003pt] =\displaystyle\sum_{j_{1}\ldots j_{m-1}}p^{(u)}_{i_{1},j_{1}\ldots j_{m-1}}q^{(v)}_{j_{1}\ldots j_{m-1},i_{2}\ldots i_{m}}\\[16.00008pt] =\displaystyle\sum_{j_{1}\ldots j_{m-1}}\Pr(X_{t+u+v}=i_{1}|X_{t+v}=j_{1},\ldots,X_{t+v-m+2}=j_{m-1})\\ \hskip 40.00006pt\cdot\Pr(X_{t+v}=j_{1},\ldots,X_{t+v-m+2}=j_{m-1}|X_{t}=i_{2},\ldots,X_{t-m+2}=i_{m})\\[8.00003pt] =\displaystyle\sum_{j_{1}\ldots j_{m-1}}\Pr(X_{t+u+v}=i_{1}|X_{t+v}=j_{1},\ldots,X_{t+v-m+2}=j_{m-1},\\[8.00003pt] \hskip 120.00018ptX_{t}=i_{2},\ldots,X_{t-m+2}=i_{m})\\[8.00003pt] \hskip 40.00006pt\cdot\Pr(X_{t+v}=j_{1},\ldots,X_{t+v-m+2}=j_{m-1}|X_{t}=i_{2},\ldots,X_{t-m+2}=i_{m})\\[8.00003pt] =\displaystyle\sum_{j_{1}\ldots j_{m-1}}\Pr(X_{t+u+v}=i_{1},X_{t+v}=j_{1},\ldots,X_{t+v-m+2}=j_{m-1}\\[8.00003pt] \hskip 70.0001pt|X_{t}=i_{2},\ldots,X_{t-m+2}=i_{m})\\[10.00002pt] =p^{(u+v)}_{i_{1}i_{2}\ldots i_{m}},\end{array}$
where in the third equality, the higher order Markovian property in (1.1) is used, while in the fourth equality, the remark below (2.2) should again be noted for the case $21\leq v\leq m-$ 2. $\Box$
3 Limiting Probability Distribution
After setting the stage in the preceding section, we are ready to review the existing lines of inquiry concerning the limiting probability distribution of the $(m-1)$th order Markov chain $X$ on state space $S=\{1,2,\ldots,n\}$ with transition tensor $\cal P$.
The first approach assumes that on the reduced first order chain $Y$, if $\lim_{t\rightarrow\infty}y_{t}$ exists, then such a limit can be approximated by a rank-one symmetric tensor
$\underbrace{y\circ\cdots\circ y}_{m-1},$ (3.1)
where $y\in\mathbb{R}^{n}$ and $\circ$ denotes the outer product. This approach can be seen in most existing works on higher order chains; see, for example, [1, 2, 3, 4, 5, 6, 8, 12, 13, 20, 21]. It also leads to the interesting $Z$-eigenvector problem. It is known [21] that such $y$ can be determined as some normalized positive $Z$-eigenvector when $\cal P$ is irreducible. As pointed out in [26], however, the approximation in (3.1) may not be able to capture the true nature of the probability distribution $y_{t}$ and its limit.
The second approach [26] involves a two-phase power iteration scheme based on the connection between the higher order chain $X$ and its reduced first order chain $Y$. It assumes that $Q$ is regular, namely $\lambda=1$ is the only dominant eigenvalue of $Q$ on the unit circle. Then, $\lim_{t\rightarrow\infty}y_{t}$ turns out to be the eigenvector of $Q$ associated with $\lambda$ which is normalized with $\|\cdot\|_{1}$. Consequently, the limiting probability distribution of $X$ follows from (2.7). As we shall show, however, that the limiting probability distribution of $X$ may still exist despite $Q$ being non-regular.
The third approach [7] deals with the higher order chain $X$ by a certain approximation scheme to its transition probabilities. The development also hinges on a direct extension of the notion of recurrent states from the first order case to the higher order case. Such an extension has been shown [11] to be an issue, which turns out to be another difference between the higher order and the first order cases.
Motivated by the above studies we have reviewed, especially by [26], our goals in this paper are to further investigate the actual limiting probability distribution of the higher order chain $X$ as well as to establish a new sufficient condition that guarantees the existence of such a limiting probability distribution. This new condition uses the regularity of $\cal P$ instead of that of $Q$ and, therefore, strengthens the existing results in [26].
We now state our main result as follows.
THEOREM 3.1
Given an $(m-1)$th order Markov chain $X=\{X_{t}:t=1,2,\ldots\}$ on state space $S$, suppose that its transition tensor $\cal P$ is regular. Then, there exists $\pi\in\mathbb{R}^{n}$ satisfying $\pi_{i}>0$ for each $i$ and $\sum_{i=1}^{n}\pi_{i}=1$ such that $\lim_{k\rightarrow\infty}p^{(k)}_{ii_{2}\ldots i_{m}}=\pi_{i}$ for all $i_{2},\ldots,i_{m}\in S$, i.e.
$\lim_{k\rightarrow\infty}{\cal P}^{(k)}=\pi\circ\underbrace{e\circ\cdots\circ e}_{m-1}.$ (3.2)
Proof: Suppose first that ${\cal P}>0$. For any $x\in\mathbb{R}^{n}$, we define the product of $x$ with any mode-$1$ fiber of ${\cal P}^{(k)}$ as
$x^{T}{\cal P}^{(k)}(:,i_{2},\ldots,i_{m})=\sum_{i=1}^{n}x_{i}p^{(k)}_{ii_{2}\ldots i_{m}}.$ (3.3)
Denote $x^{T}{\cal P}^{(k)}(:,i_{2},\ldots,i_{m})$ as $y^{(k)}_{i_{2}\ldots i_{m}}$. Clearly, $y^{(k)}=[y^{(k)}_{i_{2}\ldots i_{m}}]$ is an $(m-1)$th order, $n$-dimensional tensor, but it can also be thought of as a vector whose entries are indexed by multi-indices $i_{2}\ldots i_{m}$. By (2.1), we see
$y^{(k+1)}_{i_{2}\ldots i_{m}}=\sum_{j=1}^{n}p_{ji_{2}\ldots i_{m}}y^{(k)}_{ji_{2}\ldots i_{m-1}}.$ (3.4)
Let $\epsilon=\min_{i_{1}\ldots i_{m}}p_{i_{1}\ldots i_{m}}$, $U_{k}=\max_{i_{2}\ldots i_{m}}y^{(k)}_{i_{2}\ldots i_{m}}$, and $L_{k}=\min_{i_{2}\ldots i_{m}}y^{(k)}_{i_{2}\ldots i_{m}}$. In particular, we set $U_{0}=\max_{i}x_{i}$ and $L_{0}=\min_{i}x_{i}$. Also, without loss of generality, we assume $20<\epsilon\leq 1/$ 2.
Let $\hat{x}\in\mathbb{R}^{n}$ be obtained from $x$ by replacing all $x_{i}$ with $U_{0}$ except one entry $x_{\ell}$ that equals $L_{0}$. Since $x\leq\hat{x}$, we know that for all $i_{2},\ldots,i_{m}$,
$y^{(1)}_{i_{2}\ldots i_{m}}\leq\hat{x}^{T}{\cal P}(:,i_{2},\ldots,i_{m}).$
Concerning $\hat{x}^{T}{\cal P}(:,i_{2},\ldots,i_{m})$, we have
$\begin{split}\hat{x}^{T}{\cal P}(:,i_{2},\ldots,i_{m})&=[U_{0}e^{T}-(U_{0}-L_{0})e_{\ell}^{T}]{\cal P}(:,i_{2},\ldots,i_{m})\\[5.0pt] &=U_{0}-(U_{0}-L_{0})p_{\ell i_{2}\ldots i_{m}}\\[5.0pt] &\leq U_{0}-\epsilon(U_{0}-L_{0}),\end{split}$
which leads to $U_{1}\leq U_{0}-\epsilon(U_{0}-L_{0})$. By a similar argument in which $\hat{x}$ is obtained from $x$ by replacing all $x_{i}$ with $L_{0}$ except one entry $x_{\ell}$ which is equal to $U_{0}$, we obtain $L_{1}\geq L_{0}+\epsilon(U_{0}-L_{0})$. Thus,
$U_{1}-L_{1}\leq(1-2\epsilon)(U_{0}-L_{0}).$
From (3.4), we have $y^{(k+1)}_{i_{2}\ldots i_{m}}\leq\sum_{j=1}^{n}p_{ji_{2}\ldots i_{m}}U_{k}=U_{k}$ and so $U_{k+1}\leq U_{k}$, showing that $\{U_{k}\}$ is a non-increasing sequence. Similarly, $\{L_{k}\}$ is a non-decreasing sequence. Accordingly, $\{U_{k}-L_{k}\}$ is a non-increasing sequence. In order to prove the convergence of $\{U_{k}-L_{k}\}$, therefore, we only need to show that it has a convergent subsequence.
Applying (3.4) repeatedly, we observe that for $k>m-2$,
$y^{(k+1)}_{i_{2}\ldots i_{m}}=\sum_{j_{m-1}=1}^{n}p_{j_{m-1}i_{2}\ldots i_{m}}\sum_{j_{m-2}=1}^{n}p_{j_{m-2}j_{m-1}i_{2}\ldots i_{m-1}}\ldots\sum_{j_{1}=1}^{n}p_{j_{1}j_{2}\ldots j_{m-1}i_{2}}y^{(k-m+2)}_{j_{1}j_{2}\ldots j_{m-1}}.$
Let $\hat{y}^{(k-m+2)}$ be the vector which is obtained from $y^{(k-m+2)}$ by replacing all its entries with $U_{k-m+2}$ except one entry $y^{(k-m+2)}_{\ell_{1}\ldots\ell_{m-1}}$ which is equal to $L_{k-m+2}$. Then,
$\begin{split}y^{(k+1)}_{i_{2}\ldots i_{m}}&\leq\sum_{j_{m-1}=1}^{n}p_{j_{m-1}i_{2}\ldots i_{m}}\sum_{j_{m-2}=1}^{n}p_{j_{m-2}j_{m-1}i_{2}\ldots i_{m-1}}\ldots\sum_{j_{1}=1}^{n}p_{j_{1}j_{2}\ldots j_{m-1}i_{2}}\hat{y}^{(k-m+2)}_{j_{1}j_{2}\ldots j_{m-1}}\\ &=\sum_{j_{m-1}=1}^{n}p_{j_{m-1}i_{2}\ldots i_{m}}\sum_{j_{m-2}=1}^{n}p_{j_{m-2}j_{m-1}i_{2}\ldots i_{m-1}}\ldots\sum_{j_{1}=1}^{n}p_{j_{1}j_{2}\ldots j_{m-1}i_{2}}U_{k-m+2}\\ &\hskip 21.68121pt-p_{\ell_{m-1}i_{2}\ldots i_{m}}p_{\ell_{m-2}\ell_{m-1}i_{2}\ldots i_{m-1}}\cdots p_{\ell_{1}\ell_{2}\ldots\ell_{m-1}i_{2}}(U_{k-m+2}-L_{k-m+2})\\ \\ &\leq U_{k-m+2}-\epsilon^{m-1}(U_{k-m+2}-L_{k-m+2}),\end{split}$
which leads to
$U_{k+1}\leq U_{k-m+2}-\epsilon^{m-1}(U_{k-m+2}-L_{k-m+2}).$
Similarly, we obtain
$L_{k+1}\geq L_{k-m+2}+\epsilon^{m-1}(U_{k-m+2}-L_{k-m+2}).$
Together, these last two expressions yield
$U_{k+1}-L_{k+1}\leq(1-2\epsilon^{m-1})(U_{k-m+2}-L_{k-m+2}).$
In particular, on letting $k=m-1,2m-2,3m-3,\ldots$, we arrive at
$U_{m}-L_{m}\leq(1-2\epsilon^{m-1})(U_{1}-L_{1}),$
$U_{2m-1}-L_{2m-1}\leq(1-2\epsilon^{m-1})(U_{m}-L_{m}),$
$U_{3m-2}-L_{3m-2}\leq(1-2\epsilon^{m-1})(U_{2m-1}-L_{2m-1}),$
and so on. It follows that the subsequence
$U_{0}-L_{0},U_{1}-L_{1},U_{m}-L_{m},U_{2m-1}-L_{2m-1},U_{3m-2}-L_{3m-2},\ldots$
converges to zero, and hence so does the entire sequence $\{U_{k}-L_{k}\}$.
The above analysis also shows that
$\{U_{0},U_{1},U_{m},U_{2m-1},U_{3m-2},\ldots\}$
is strictly decreasing while
$\{L_{0},L_{1},L_{m},L_{2m-1},L_{3m-2},\ldots\}$
is strictly increasing unless $U_{k}-L_{k}=0$ for some $k=0,1,m,2m-1,3m-2,\ldots$. In particular, if we start with $U_{0}-L_{0}>0$, then there must be some $L_{0}<c<U_{0}$ such that $\lim_{k\rightarrow\infty}(U_{k}-L_{k})=0$, i.e. $\lim_{k\rightarrow\infty}y^{(k)}_{i_{2}\ldots i_{m}}=c$ for all $i_{2},\ldots,i_{m}$.
Next, let us denote by $e_{i}$, $1\leq i\leq n$, the $i$th column of an $n\times n$ identity matrix. For each $i$, we pick $x=e_{i}$ in (3.3) and apply the above convergence result. In this case, $L_{0}=0$ and $U_{0}=1$. It follows, therefore, that for each $i$, there exists $10<\pi_{i}<$ 1 such that $\lim_{k\rightarrow\infty}p^{(k)}_{ii_{2}\ldots i_{m}}=\pi_{i}$ for all $i_{2},\ldots,i_{m}$. The fact that $\sum_{i=1}^{n}\pi_{i}=1$ is obvious because of the stochasticity of ${\cal P}^{(k)}$.
Finally, in the case that $\cal P$ is regular but ${\cal P}\not>0$, there exists $K\geq 1$ such that ${\cal P}^{(K)}>0$. We set $\epsilon=\min_{i_{1}\ldots i_{m}}p^{(K)}_{i_{1}\ldots i_{m}}$ and consider the subsequence
$U_{0}-L_{0},U_{K}-L_{K},U_{mK}-L_{mK},U_{(2m-1)K}-L_{(2m-1)K},U_{(3m-2)K}-L_{(3m-2)K},\ldots.$
The proof can be done in exactly the same fashion as the ${\cal P}>0$ case. $\Box$
For the case of second order chains, a result similar to Theorem 3.1 can be found in [25].
Our next result confirms that $\pi$ in Theorem 3.1 is indeed the limiting probability distribution of the chain $X$ which contains the long-term probabilities for the chain to be in various states. Moreover, these probabilities are independent of the initial distribution of $X_{1},X_{2},\ldots,X_{m-1}$.
THEOREM 3.2
Suppose that $X=\{X_{t}:t=1,2,\ldots\}$ is an $(m-1)$th order Markov chain on state space $S$ whose transition tensor is $\cal P$. If there exists a nonnegative $\pi\in\mathbb{R}^{n}$ such that $\sum_{i=1}^{n}\pi_{i}=1$ and ${\lim_{k\rightarrow\infty}{\cal P}^{(k)}=\pi\circ\underbrace{e\circ\cdots\circ e}_{m-1}}$, then
$\lim_{t\rightarrow\infty}\Pr(X_{t}=i)=\pi_{i}$
for each $i\in S$.
Proof: For any $k\geq 1$, we observe
$\begin{split}\Pr(X_{m+k-1}=i)&=\sum_{i_{2},\ldots,i_{m}=1}^{n}\Pr(X_{m+k-1}=i,X_{m-1}=i_{2},\ldots,X_{1}=i_{m})\\[5.0pt] &=\sum_{i_{2},\ldots,i_{m}=1}^{n}\Pr(X_{m+k-1}=i|X_{m-1}=i_{2},\ldots,X_{1}=i_{m})\\ &\hskip 57.81621pt\cdot\Pr(X_{m-1}=i_{2},\ldots,X_{1}=i_{m})\\[5.0pt] &=\sum_{i_{2},\ldots,i_{m}=1}^{n}p^{(k)}_{ii_{2}\ldots i_{m}}\Pr(X_{m-1}=i_{2},\ldots,X_{1}=i_{m}).\end{split}$
Pushing $k\rightarrow\infty$ and noting $\sum_{i_{2},\ldots,i_{m}=1}^{n}\Pr(X_{m-1}=i_{2},\ldots,X_{1}=i_{m})=1$, the proof is complete. $\Box$
We comment here that Theorem 3.2 does not require $\pi$ to be positive. This will allow us to explore in future works limiting behavior of the chain $X$ in broader settings such as higher order absorbing chains.
From Theorems 3.1 and 3.2, we see that the limiting probability distribution of the chain $X$ can be investigated entirely within a tensor theoretic framework without resorting to the reduced first order chain or relying on the two-phase power iteration scheme as in [26].
Theorem 3.1 establishes a sufficient condition, that is, $\cal P$ being regular, for the existence of the limiting probability distribution of the chain $X$. In addition, it shows that in this case, the limiting probability distribution is positive. This extends the well-known result regarding first order chains as recalled in the introductory section.
Besides, we mention that without the condition of $\cal P$ being regular, the limiting probability distribution may not exist — see the following example. Hence, such a sufficient condition is “tight”.
EXAMPLE 3.1
Let us consider a second order chain on the state space $\{1,2,3\}$ with transition tensor
${\cal P}(:,:,i_{3})=\left[\begin{array}[]{ccc}0&1/2&0\\ 1&0&1\\ 0&1/2&0\end{array}\right]$
for any $i_{3}=1,2,3$. This $\cal P$ is ergodic but is non-regular. When $k$ is odd, we have ${\cal P}^{(k)}={\cal P}$. When $k$ is even, we have
${\cal P}^{(k)}(:,:,i_{3})=\left[\begin{array}[]{ccc}1/2&0&1/2\\ 0&1&0\\ 1/2&0&1/2\end{array}\right]$
for $i_{3}=1,2,3$. As a result, $\lim_{k\rightarrow\infty}{\cal P}^{(k)}$ does not exist.
To justify that this chain has no limiting probability distribution, let us assume $X_{1}=X_{2}=1$. Then, $x_{t}$ may be determined by (2.5). Alternatively, we can proceed as follows.
By $p_{211}=1$, we see $X_{3}=2$, i.e. $x_{3}=[0~{}1~{}0]^{T}$. Next, using $p_{121}=p_{321}=1/2$, the chain will move to either state $1$ or state $3$ with equal probability, i.e. $x_{4}=[1/2~{}~{}0~{}~{}1/2]^{T}$. Continuing, since $p_{212}=p_{232}=1$, the chain will return to state $2$ in its next move, i.e. $x_{5}=[0~{}1~{}0]^{T}$. Hence, $x_{t}$ will bounce back and forth between $[0~{}1~{}0]^{T}$ and $[1/2~{}~{}0~{}~{}1/2]^{T}$ and will not converge as $t\rightarrow\infty$.
From the proof of Theorem 3.2, we can also see the role that the limit ${\lim_{k\rightarrow\infty}{\cal P}^{(k)}=\pi\circ\underbrace{e\circ\cdots\circ e}_{m-1}}$ plays so as to guarantee that the limiting probability distribution is independent of the initial probability distribution of $X_{1},X_{2},\ldots,X_{m-1}$. This result can be extended slightly as follows.
THEOREM 3.3
Suppose that $X=\{X_{t}:t=1,2,\ldots\}$ is an $(m-1)$th order Markov chain on state space $S$ and with transition tensor $\cal P$. If there exists some tensor $\cal R$ of the same size as $\cal P$ such that $\lim_{k\rightarrow\infty}{\cal P}^{(k)}=\cal R$, then for each $i\in S$, $\lim_{t\rightarrow\infty}\Pr(X_{t}=i)$ exists but, in general, it depends on the initial probability distribution of $X_{1},X_{2},\ldots,X_{m-1}$.
Proof: Similar to the proof of Theorem 3.2, for any $k\geq 1$, we find
$\Pr(X_{m+k-1}=i)=\sum_{i_{2},\ldots,i_{m}=1}^{n}p^{(k)}_{ii_{2}\ldots i_{m}}\Pr(X_{m-1}=i_{2},\ldots,X_{1}=i_{m}).$
On letting ${\cal R}=[r_{i_{1}i_{2}\ldots i_{m}}]$, it follows that
$\lim_{t\rightarrow\infty}\Pr(X_{t}=i)=\sum_{i_{2},\ldots,i_{m}=1}^{n}r_{ii_{2}\ldots i_{m}}\Pr(X_{m-1}=i_{2},\ldots,X_{1}=i_{m}).$ (3.5)
$\Box$
In view of Theorem 3.3, we now revisit Example 2.1.
EXAMPLE 3.2
Given the chain in Example 2.1, it is easy to verify that
$\lim_{k\rightarrow\infty}{\cal P}^{(k)}(:,:,1)=\left[\begin{array}[]{cc}1&1/3\\ 0&2/3\end{array}\right]~{}{\rm and}~{}\lim_{k\rightarrow\infty}{\cal P}^{(k)}(:,:,2)=\left[\begin{array}[]{cc}2/3&0\\ 1/3&1\end{array}\right].$
According to (3.5), we obtain that the entries of $\lim_{t\rightarrow\infty}x_{t}$ are given by
$\lim_{t\rightarrow\infty}\Pr(X_{t}=1)=\Pr(X_{2}=1,X_{1}=1)+\frac{\Pr(X_{2}=2,X_{1}=1)}{3}+\frac{2\Pr(X_{2}=1,X_{1}=2)}{3}$
and
$\lim_{t\rightarrow\infty}\Pr(X_{t}=2)=\frac{2\Pr(X_{2}=2,X_{1}=1)}{3}+\frac{\Pr(X_{2}=1,X_{1}=2)}{3}+\Pr(X_{2}=2,X_{1}=2).$
Our final result connects the limiting probability distribution with the dominant eigenvectors of $Q$. A similar result for a regular $Q$ is provided in [26] too.
THEOREM 3.4
Let $\pi$ be the limiting probability distribution of an $(m-1)$th order regular Markov chain $X$ on state space $S$. Let $Q$ be the transition matrix of its reduced first order chain $Y$ and assume that $z\in\mathbb{R}^{N}$ is a normalized nonnegative right eigenvector of $Q$ associated with dominant eigenvalue $1$, i.e. $Qz=z$, $z\geq 0$, and $e^{T}z=1$. Then,
$\pi=P^{(0)}z,$ (3.6)
where $P^{(0)}$ is given in (2.3).
Proof: According to (2.7), we have
$P^{(k)}z=P^{(0)}Q^{k}z=P^{(0)}z.$
From (3.2), we clearly have $\lim_{k\rightarrow\infty}P^{(k)}=\pi e^{T}$. Hence
$\pi=P^{(0)}z,$
i.e. (3.6) holds. $\Box$
We remark that the above result holds even when $Q$ is non-regular and has multiple dominant eigenvectors associated with eigenvalues of modulus $1$. In fact, this may happen even when $\cal P$ is regular, as illustrated by the example below from [7].
EXAMPLE 3.3
Take a second order chain on the state space $\{1,2,3,4\}$ whose transition tensor $\cal P$ has the form
${\cal P}(:,:,1)=\left[\begin{array}[]{cccc}1/2&0&0&0\\ 1/2&0&1&0\\ 0&1&0&1\\ 0&0&0&0\end{array}\right],~{}{\cal P}(:,:,2)=\left[\begin{array}[]{cccc}0&0&1/2&1\\ 0&1/2&0&0\\ 1/2&1/2&0&0\\ 1/2&0&1/2&0\end{array}\right],$
${\cal P}(:,:,3)=\left[\begin{array}[]{cccc}0&1&0&1\\ 1&0&1/2&0\\ 0&0&1/2&0\\ 0&0&0&0\end{array}\right],~{}{\cal P}(:,:,4)=\left[\begin{array}[]{cccc}0&0&0&0\\ 1&1&1&0\\ 0&0&0&1/2\\ 0&0&0&1/2\end{array}\right].$
Using (2.1), we obtain ${\cal P}^{(10)}>0$ and thus $\cal P$ is regular. Theorem 3.1 guarantees the existence of the limiting probability distribution $\pi$ of this chain. On the other hand, with (2.4), it is easy to verify that the transition matrix $Q$ of the reduced first order chain has a double dominant eigenvalue $\lambda=1$ with an eigenspace spanned by
$z^{(1)}=\frac{1}{7}[0~{}0~{}1~{}1~{}2~{}0~{}0~{}0~{}0~{}2~{}0~{}0~{}0~{}0~{}1~{}0]^{T}$
and
$z^{(2)}=\frac{1}{7}[0~{}2~{}0~{}0~{}0~{}0~{}2~{}0~{}1~{}0~{}0~{}1~{}1~{}0~{}0~{}0]^{T}.$
Because of this, the two-phase power iteration method in [26] is no longer appropriate since it does not converge to a unique answer. Nevertheless, it follows from (3.6) that
$\pi=P^{(0)}z^{(1)}=P^{(0)}z^{(2)}=\frac{1}{7}[2~{}2~{}2~{}1]^{T}.$